OpenStudy (anonymous):

What does a common solution mean in algebra 2?? please help?

OpenStudy (anonymous):

@Calcmathlete can you please help me??

OpenStudy (anonymous):

Well, what's the problem this is from?

OpenStudy (anonymous):

Otherwise though, a common solution means same solution.

OpenStudy (anonymous):

THis also means that they share the solution.

OpenStudy (anonymous):

It varies on the problem and the context.

OpenStudy (anonymous):

can i show you the problem??

OpenStudy (anonymous):

Which pair of equations does NOT have a common solution? A) x + y = -14x - 3y = 24 B) 4x + 6y = 12 6x + 9y = 12 C) 2x - 3y = -4 2x + y = 4 D) 5x - 4y = 9 x - 2y = -3.. if you can solve this an explain your answer this would BE VERYYYYY HELPFUL

OpenStudy (anonymous):

Are you familiar with the cases of no solution and infinitely many solutions, 0 = 0, 0 = 6 respectively?

OpenStudy (anonymous):

a little bit.. don't remember it...

OpenStudy (anonymous):

Well, in some cases, the variable will disappear. If you are left with a situation like this: 0 = 0 or 6 = 6, then you have infinitely many solutions meaning that it's the sme line. If you have a case that's like 0 = 4 or 5 = 7, then you have a case of no solution in which they re parallel lines and therefore never intersect. Set up a system of equations for each choice and see which one comes out like 0 = 4 or 5 = 2, etc.

OpenStudy (anonymous):

oh so i this case it has to be a common solution would have to be 0=0? ???

OpenStudy (anonymous):

We're looking for the case in which it's NO SOLUTION. this means that the variable has to disappear and the numbers have to be different. See how the numbers are different?: 4 = 7

OpenStudy (anonymous):

ohhhhh yeah! okay thanks :) i see now,. you explained that great!! :D

OpenStudy (anonymous):

No Problem :)

OpenStudy (anonymous):

-7/8=8/7 does that give us our answer.. thats what i ended up with on the first one A...